There are two categories of ultraviolet light. Ultraviolet A ( UVA ) has a wavelength ranging from 320nm to 400nm . It is not so harmful to the skin and is necessary for the production of vitamin D. UVB, with a wavelength between 280nm and 320nm , is much more dangerous, because it causes skin cancer

PART A
Find the frequency ranges of UVA.
Enter your answers separated with commas.

fmin,fmax= 7.50
Answer #1

a)f=c/λ

Range=3*108/400*10-9 to 3*108/320*10-9

=7.5*1014,9.375*1014

b)Range=9.375*1014,1.07*1015

c)Wave number=2π/λ

Range=1.57*107,1.96*107

d)Range=1.96*107,2.24*107